# Modified Langevin noise formalism for multiple quantum emitters in dispersive electromagnetic environments

### v1

@@ Appendix A Bright electric field operator
$\hat{\mathbf{E}}^{bright}(\mathbf{r})=\sum_{j=1}^{N}\int_{0}^{\infty}d\omega\left[\mathbf{E}_{j}^{(M)}(\mathbf{r},\omega)\hat{C}_{j}(\omega)+h.c.\right]+\\ \sum_{j=1}^{N}\int_{0}^{\infty}d\omega\left[\mathbf{E}_{j}^{(S)}(\mathbf{r},\omega)\hat{D}_{j}(\omega)+h.c.\right]$ (76)
@@
$\mathbf{E}_{j}^{(M)}(\mathbf{r},\omega)$ $=\sum_{j=1}^{N}U_{ij}^{*}(\omega)\mathbf{e}_{j}^{(M)}(\mathbf{r},\omega),$ (77a)
$\mathbf{E}_{j}^{(S)}(\mathbf{r},\omega)$ $=\sum_{j=1}^{N}V_{ij}^{*}(\omega)\mathbf{e}_{j}^{(S)}(\mathbf{r},\omega),$ (77b)
@@
$\mathbf{e}_{j}^{(M)}(\mathbf{r},\omega)=\int_{V}d^{3}\mathbf{r}^{\prime}\mathcal{G}_{eq}(\mathbf{r},\mathbf{r}^{\prime},\omega)\mathcal{G}^{*T}_{eq}(\mathbf{r}^{\prime},\mathbf{r}_{j},\omega)\mathbb{u}_{j},$ (78a)
$\mathbf{e}_{j}^{(S)}(\mathbf{r},\omega)=\int do_{\mathbb{n}}\sum_{\nu}\mathbf{E}_{\omega\mathbb{n}\nu}(\mathbf{r})\mathbf{E}_{\omega\mathbb{n}\nu}(\mathbf{r}_{j})\mathbb{u}_{j}.$ (78b)
@@ Appendix B Spectral density matrices and power observables in the framework of classical electrodynamics
We here show that the spectral density matrices, introduced in Section IV, are related to power observables of the system of multiple emitters and dielectric objects in the framework of classical electrodynamics. We consider a reciprocal dielectric object in an unbounded space. The object is driven by $N$ electric dipoles, all oscillating harmonically at the angular frequency $\omega$. The $i$-the electric dipole is located at position $\mathbf{r}_{i}$ and has dipole moment $\mathbf{p}_{i}=\mu_{i}\mathbf{u}_{i}$ where $\mu_{i}$, assumed real, is the amplitude and $\mathbf{u}_{i}$ is a unit vector that gives the dipole orientation. The associated electric current density is $\mathbf{J}_{i}=-i\omega\mu_{i}\mathbf{u}_{i}\,\delta\,(\mathbf{r}-\mathbf{r}_{i})$. Because the medium is linear, the total electric field generated by $\mathbf{J}(\mathbf{r})=\sum_{i=1}^{N}\mathbf{J}_{i}(\mathbf{r})$ is the superposition of the fields produced when each dipole acts alone:
@@ B.1 Lorentz Reciprocity and a far-field identity
$\mathbf{E}_{i}(\mathbf{r}_{\infty}){\approx}\frac{\mathbf{A}_{i}(\mathbf{n})}{r_{\infty}}\,e^{ik_{\omega}r_{\infty}}$ (83)
@@
$\mathbf{E}_{\infty}^{inc}(\mathbf{r}_{i})\approx\frac{k_{\omega}^{2}}{4\pi\varepsilon_{0}}\,\frac{\mu_{\infty}}{r_{\infty}}\exp\left(ik_{\omega}\,\mathbf{n}\cdot\mathbf{r}_{i}\right)\exp\left(-ik_{\omega}r_{\infty}\right)\mathbf{u}_{\infty}$ (84)
as $\mathbf{r}_{i}$ lies in the Fraunhofer zone of the electromagnetic field generated by the dipole $\mathbb{p}_{\infty}$. Thus, in Scenario II, the dielectric object experiences an incident field that is a transverse plane wave propagating along $\mathbf{n}$, polarized along $\mathbf{u}_{\infty}$ and with unit amplitude provided that $\mu_{\infty}=4\pi\varepsilon_{0}\,r_{\infty}/{k_{\omega}^{2}}$. According to Eqs. 7-8, we denote by $\mathbf{F}_{\omega\mathbf{n}\mathbf{u}_{\infty}}(\mathbf{r}_{i})$ the total electric field at $\mathbf{r}_{i}$ scattered by the dielectric objects when excited by a plane wave of unit intensity propagating along $\mathbf{n}$ and polarized along $\mathbf{u}_{\infty}$, hence
@@
This relation states that the $\mathbf{u}_{\infty}$-polarized far-field amplitude in direction $\mathbf{n}$ generated by the dipole at $\mathbb{p}_{i}$ is proportional to the projection of the normalized scattering mode $\mathbf{E}_{\omega\mathbf{n}\mathbf{u}_{\infty}}$ at the dipole location onto the dipole orientation $\mathbf{u}_{i}$. In other words, it links the coupling between the dipole and the scattering mode at $\mathbf{r}_{i}$ to the corresponding far-field pattern produced when the object is driven by a dipole $\mu_{i}\mathbf{u}_{i}$ at $\mathbf{r}_{i}$.
@@ B.2 Time-averaged power emitted by the dipoles
Using 80 in 89, one finds the following
@@ B.3 Time-averaged power absorbed by the dielectric object
$\mathscr{P}^{abs}=\frac{\omega}{2}\varepsilon_{0}\int_{V}d^{3}\mathbf{r}\,\text{Im}[\varepsilon_{\omega}(\mathbf{r})]\mathbf{E}(\mathbf{r})\cdot\mathbf{E}^{*}(\mathbf{r})\\ =\sum_{i,j=1}^{N}\mathscr{P}^{abs}_{ij}$ (92)
@@
Using 80 and the reciprocity for the dyadic Green function we get:
$\mathscr{P}^{abs}_{ij}=\frac{1}{2}\varepsilon_{0}\mu_{0}^{2}\omega^{5}\mu_{i}\mu_{j}\times\\ \int_{V}d^{3}\text{Im}[\varepsilon_{\omega}(\mathbf{r})]\mathbf{u}_{i}\cdot\mathcal{G}_{\omega}(\mathbf{r}_{i},\mathbf{r})\mathcal{G}_{\omega}^{*T}(\mathbf{r}_{j},\mathbf{r})\cdot\mathbf{u}_{j}$ (94)
@@
$\mathscr{P}^{abs}_{ij}=\frac{\pi}{2}{\hbar\omega}\frac{\mu_{i}\mu_{j}}{\hbar^{2}}\times\\ \int_{V}d^{3}\mathbf{r}\,\mathbf{u}_{i}\cdot\mathcal{G}_{e}(\mathbf{r}_{i},\mathbf{r},\omega)\mathcal{G}_{e}^{*T}(\mathbf{r}_{j},\mathbf{r},\omega)\cdot\mathbf{u}_{j}=\frac{\pi}{2}{\hbar\omega}\mathcal{J}^{(M)}_{ij}$ (95)
@@ B.4 Averaged Power Radiated to infinity
$\mathscr{P}^{rad}=\frac{1}{4}\int_{S_{\infty}}\,do_{\mathbf{n}}r_{\infty}^{2}\left(\mathbf{E}\times\mathbf{H}^{*}+\mathbf{E}^{*}\times\mathbf{H}\right)\cdot\mathbf{n}=\\ \sum_{i,j=1}^{N}\mathscr{P}^{rad}_{ij}$ (96)
@@
The Poynting vector 98 becomes
@@
By replacing this expression in 97 we get:
$\mathscr{P}^{rad}_{ij}=\frac{1}{2\zeta_{0}}\oint_{S_{\infty}}\,do_{\mathbf{n}}\,\displaystyle\sum_{\nu}(\mathbf{A}_{i}\cdot\bm{\nu})(\mathbf{A}_{j}^{*}\cdot\bm{\nu})$ (101)
Using relation 87 one finds
$\mathscr{P}^{rad}_{ij}=\frac{\pi}{2}{\hbar\omega}\frac{\mu_{i}\mu_{j}}{\hbar^{2}}\times\\ \oint_{S_{\infty}}\,do_{\mathbf{n}}\,\sum_{\nu}\mathbf{u}_{i}\cdot\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r}_{i})\mathbf{E}^{*}_{\omega\mathbf{n}\nu}(\mathbf{r}_{j})\cdot\mathbf{u}_{j}=\frac{\pi}{2}{\hbar\omega}\mathrm{\mathcal{J}}_{ij}^{(S)}(\omega)$ (102)
@@ B.5 Mutual Poynting theorem
$\mathscr{P}_{ij}^{abs}=\frac{\omega}{2}\varepsilon_{0}\int_{V}d^{3}\mathbf{r}\,\text{Im}[\varepsilon_{\omega}(\mathbf{r})]\mathbf{E}_{i}(\mathbf{r})\cdot\mathbf{E}_{j}^{*}(\mathbf{r})\\ =-\oint_{\partial\Omega}d^{2}\mathbf{r}\,\mathbf{S}_{ij}\,\cdot\mathbf{m}$ (103)
@@ Appendix C Correlators
In this Section we compute the elements $C_{ij}^{(M)}(t)$ of the correlator matrix of the medium-assisted reservoir; a similar result holds for the elements of the correlator matrix of the scattering-assisted reservoir.
We consider the following scenario: (a) the initial quantum state of the electromagnetic environment is a product state, that is, $\hat{\rho}_{E}(0)=\hat{\rho}_{E}^{(M)}(0)\otimes\hat{\rho}_{E}^{(S)}$, where $\hat{\rho}_{E}^{(M)}(0)$ and $\hat{\rho}_{E}^{(S)}(0)$ are the initial density operators of the medium-assisted reservoir and the scattering-assisted reservoir, respectively; (b) the two reservoirs are initially in thermal states with temperatures $T_{0}^{(M)}$ and $T_{0}^{(S)}$.
The correlation function $C^{(M)}_{ij}(t)$ is given by
$C^{(M)}_{ij}(t)=\langle\hat{F}^{(M)}_{i}(t)\hat{F}^{(M)}_{j}(0)\rangle$ (104)
where $\hat{{F}}^{(M)}_{i}(t)$ is the interaction operator of the medium-assisted reservoir in the interaction picture,
$\hat{{F}}^{(M)}_{i}(t)=\hat{U}^{\dagger}_{E}(t)\hat{F}^{(M)}_{i}\hat{U}_{E}(t),$ (105)
$\hat{U}_{E}(t)=\exp(-i\hat{H}_{E}^{bright}t/\hbar)$ is the free evolution operator of the bright bosonic modes of the electromagnetic environment, and $\langle\cdot\rangle=Tr_{E_{M}}[\cdot\,\rho_{E}^{(M)}(0)]$. Starting with 105, and using Eqs. 32, 46, and the commutation relations 30 we obtain
$\hat{F}_{i}^{(M)}(t)=\sum_{j=1}^{N}\int_{0}^{\infty}d\omega[G_{ij}^{(M)}(\omega)\hat{C}_{j}(\omega)e^{-i\omega t}+h.c.].$ (106)
Substituting 106 into 104, and using again 30 we find
$C_{ij}^{(M)}(t)=\int_{0}^{\infty}d\omega\sum_{h=1}^{N}\left[G^{(M)}_{in}(\omega)G^{(M)*}_{jn}(\omega)\langle\hat{C}_{n}(\omega)\hat{C}^{\dagger}_{n}(\omega)\rangle e^{-i\omega t}+G^{(M)*}_{in}(\omega)G^{(M)}_{jn}(\omega)\langle\hat{C}^{\dagger}_{n}(\omega)\hat{C}_{n}(\omega)\rangle e^{+i\omega t}\right],$ (107)
@@
$C_{ij}^{(M)}(t)=\int_{0}^{\infty}\mathrm{d}\omega[(1+n_{\omega}^{(M)}){M}_{ij}(\omega)e^{-i\omega t}+n_{\omega}^{(M)}{M}_{ij}^{*}(\omega)e^{i\omega t}].$ (110)
@@ Appendix D Surface Integral Equation Method
Following Poggio–Miller–Chang–Harrington–Wu–Tsai (PMCHWT) [62, 63, 64], the equivalent electric and magnetic surface currents $\mathbf{j}_{e}(\mathbf{r})$ and $\mathbf{j}_{m}(\mathbf{r})$, defined on $\partial V$, solve:
$\bm{\mathcal{Z}}\,\mathbf{j}=\mathbf{v},$ (111)
@@
$\mathbf{j}=\left[\mathbf{j}_{e},\mathbf{j}_{m}\right]^{T}$, $\mathbf{v}=\left[\mathbf{e}_{0},\mathbf{h}_{0}\right]^{T}$, $\mathbf{e}_{0}=\left.-\mathbf{n}\times\mathbf{n}\times{\bf{E}}_{inc}\right|_{\partial\Omega}$, and $\mathbf{h}_{0}=\left.-\mathbf{n}\times\mathbf{n}\times{\bf{H}}_{inc}\right|_{\partial\Omega}.$ The EFIE/MFIE boundary operators $\mathcal{T}_{\pm}$ and $\mathcal{K}_{\pm}$ act on a tangential test function $\mathbf{w}$ as:
$\bm{\mathcal{K}}_{\pm}$ $\left\{{\bf w}\right\}\left({\bf r}\right)=\mathbf{n}\times\mathbf{n}\times\int_{\partial\Omega}{\bf w}\left({\bf r}^{\prime}\right)\times\nabla^{\prime}g^{\pm}\left({\bf r}-{\bf r}^{\prime}\right)dS^{\prime},$ (113a)
$\bm{\mathcal{T}}_{\pm}$ $\left\{{\bf w}\right\}\left({\bf r}\right)=-ik^{\pm}_{\omega}\mathbf{n}\times\mathbf{n}\times\int_{\partial\Omega}g^{\pm}\left({\bf r}-{\bf r}^{\prime}\right){\bf w}\left({\bf r}^{\prime}\right)dS^{\prime}$
$-\frac{1}{ik^{\pm}_{\omega}}\mathbf{n}\times\mathbf{n}\times\int_{\partial\Omega}\nabla^{\prime}g^{\pm}\left({\bf r}-{\bf r}^{\prime}\right)\nabla^{\prime}_{S}\cdot{\bf w}\left({\bf r}^{\prime}\right)dS^{\prime},$ (113b)

### v2

@@ Appendix A Bright electric field operator
$\hat{\mathbf{E}}^{\text{bright}}(\mathbf{r})=\sum_{j=1}^{N}\int_{0}^{\infty}d\omega\left[\mathbf{E}_{j}^{(M)}(\mathbf{r},\omega)\hat{C}_{j}(\omega)+h.c.\right]+\\ \sum_{j=1}^{N}\int_{0}^{\infty}d\omega\left[\mathbf{E}_{j}^{(S)}(\mathbf{r},\omega)\hat{D}_{j}(\omega)+h.c.\right],$ (74)
@@
$\mathbf{E}_{i}^{(M)}(\mathbf{r},\omega)$ $=\sum_{j=1}^{N}U_{ij}^{*}(\omega)\mathbf{e}_{j}^{(M)}(\mathbf{r},\omega),$ (75a)
$\mathbf{E}_{i}^{(S)}(\mathbf{r},\omega)$ $=\sum_{j=1}^{N}V_{ij}^{*}(\omega)\mathbf{e}_{j}^{(S)}(\mathbf{r},\omega),$ (75b)
@@
$\mathbf{e}_{j}^{(M)}(\mathbf{r},\omega)=\int_{V}d^{3}\mathbf{r}^{\prime}\mathcal{G}_{e}(\mathbf{r},\mathbf{r}^{\prime},\omega)\mathcal{G}^{*T}_{e}(\mathbf{r}^{\prime},\mathbf{r}_{j},\omega)\mathbf{u}_{j},$ (76a)
$\mathbf{e}_{j}^{(S)}(\mathbf{r},\omega)=\int do_{\mathbf{n}}\sum_{\nu}\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r})\mathbf{E}_{\omega\mathbf{n}\nu}^{*}(\mathbf{r}_{j})\mathbf{u}_{j}.$ (76b)
@@ Appendix B Spectral density matrices and power observables in the framework of classical electrodynamics
We here show that the spectral density matrices, introduced in Section IV, are related to power observables of the system of multiple emitters and dielectric objects in the framework of classical electrodynamics. We consider a reciprocal dielectric object in an unbounded space. The object is driven by $N$ electric dipoles, all oscillating harmonically at the angular frequency $\omega$. The $i$-th electric dipole is located at position $\mathbf{r}_{i}$ and has dipole moment $\mathbf{p}_{i}=\mu_{i}\mathbf{u}_{i}$ where $\mu_{i}$, assumed real, is the amplitude and $\mathbf{u}_{i}$ is a unit vector that gives the dipole orientation. The associated electric current density is $\mathbf{J}_{i}=-i\omega\mu_{i}\mathbf{u}_{i}\,\delta\,(\mathbf{r}-\mathbf{r}_{i})$. Because the medium is linear, the total electric field generated by $\mathbf{J}(\mathbf{r})=\sum_{i=1}^{N}\mathbf{J}_{i}(\mathbf{r})$ is the superposition of the fields produced when each dipole acts alone:
@@ B.1 Lorentz Reciprocity and a far-field identity
$\mathbf{E}_{i}(\mathbf{r}_{\infty}){\approx}\frac{\mathbf{A}_{i}(\mathbf{n})}{r_{\infty}}\,e^{ik_{\omega}r_{\infty}},$ (81)
@@
$\mathbf{E}_{\infty}^{inc}(\mathbf{r}_{i})\approx\frac{k_{\omega}^{2}}{4\pi\varepsilon_{0}}\,\frac{\mu_{\infty}}{r_{\infty}}\exp\left(ik_{\omega}\,\mathbf{n}\cdot\mathbf{r}_{i}\right)\exp\left(-ik_{\omega}r_{\infty}\right)\mathbf{u}_{\infty},$ (82)
as $\mathbf{r}_{i}$ lies in the Fraunhofer zone of the electromagnetic field generated by the dipole $\mathbf{p}_{\infty}$. Thus, in Scenario II, the dielectric object experiences an incident field that is a transverse plane wave propagating along $\mathbf{n}$, polarized along $\mathbf{u}_{\infty}$ and with unit amplitude provided that $\mu_{\infty}=4\pi\varepsilon_{0}\,r_{\infty}/{k_{\omega}^{2}}$. According to Eqs. 7-8, we denote by $\mathbf{F}_{\omega\mathbf{n}\mathbf{u}_{\infty}}(\mathbf{r}_{i})$ the total electric field at $\mathbf{r}_{i}$ scattered by the dielectric objects when excited by a plane wave of unit intensity propagating along $\mathbf{n}$ and polarized along $\mathbf{u}_{\infty}$, hence
@@
This relation states that the $\mathbf{u}_{\infty}$-polarized far-field amplitude in direction $\mathbf{n}$ generated by the dipole at $\mathbf{p}_{i}$ is proportional to the projection of the normalized scattering mode $\mathbf{E}_{\omega\mathbf{n}\mathbf{u}_{\infty}}$ at the dipole location onto the dipole orientation $\mathbf{u}_{i}$. In other words, it links the coupling between the dipole and the scattering mode at $\mathbf{r}_{i}$ to the corresponding far-field pattern produced when the object is driven by a dipole $\mu_{i}\mathbf{u}_{i}$ at $\mathbf{r}_{i}$.
@@ B.2 Time-averaged power emitted by the dipoles
Using 78 in 87, one finds the following
@@ B.3 Time-averaged power absorbed by the dielectric object
$\mathscr{P}^{abs}=\frac{\omega}{2}\varepsilon_{0}\int_{V}d^{3}\mathbf{r}\,\text{Im}[\varepsilon_{\omega}(\mathbf{r})]\mathbf{E}(\mathbf{r})\cdot\mathbf{E}^{*}(\mathbf{r})\\ =\sum_{i,j=1}^{N}\mathscr{P}^{abs}_{ij},$ (90)
@@
Using 78 and the reciprocity for the dyadic Green function, we get:
$\mathscr{P}^{abs}_{ij}=\frac{1}{2}\varepsilon_{0}\mu_{0}^{2}\omega^{5}\mu_{i}\mu_{j}\times\\ \int_{V}d^{3}\text{Im}[\varepsilon_{\omega}(\mathbf{r})]\mathbf{u}_{i}\cdot\mathcal{G}_{\omega}(\mathbf{r}_{i},\mathbf{r})\mathcal{G}_{\omega}^{*T}(\mathbf{r}_{j},\mathbf{r})\cdot\mathbf{u}_{j},$ (92)
@@
$\mathscr{P}^{abs}_{ij}=\frac{\pi}{2}{\hbar\omega}\frac{\mu_{i}\mu_{j}}{\hbar^{2}}\times\\ \int_{V}d^{3}\mathbf{r}\,\mathbf{u}_{i}\cdot\mathcal{G}_{e}(\mathbf{r}_{i},\mathbf{r},\omega)\mathcal{G}_{e}^{*T}(\mathbf{r}_{j},\mathbf{r},\omega)\cdot\mathbf{u}_{j}=\frac{\pi}{2}{\hbar\omega}\mathcal{J}^{(M)}_{ij}$, (93)
@@ B.4 Averaged Power Radiated to infinity
$\mathscr{P}^{rad}=\frac{1}{4}\int_{S_{\infty}}\,do_{\mathbf{n}}r_{\infty}^{2}\left(\mathbf{E}\times\mathbf{H}^{*}+\mathbf{E}^{*}\times\mathbf{H}\right)\cdot\mathbf{n}=\\ \sum_{i,j=1}^{N}\mathscr{P}^{rad}_{ij},$ (94)
@@
The Poynting vector 96 becomes
@@
By replacing this expression in 95 we get:
$\mathscr{P}^{rad}_{ij}=\frac{1}{2\zeta_{0}}\oint_{S_{\infty}}\,do_{\mathbf{n}}\,\displaystyle\sum_{\nu}(\mathbf{A}_{i}\cdot\boldsymbol{\nu})(\mathbf{A}_{j}^{*}\cdot\boldsymbol{\nu}).$ (99)
Using the relation 85 one finds
$\mathscr{P}^{rad}_{ij}=\frac{\pi}{2}{\hbar\omega}\frac{\mu_{i}\mu_{j}}{\hbar^{2}}\times\\ \oint_{S_{\infty}}\,do_{\mathbf{n}}\,\sum_{\nu}\mathbf{u}_{i}\cdot\mathbf{E}_{\omega\mathbf{n}\nu}(\mathbf{r}_{i})\mathbf{E}^{*}_{\omega\mathbf{n}\nu}(\mathbf{r}_{j})\cdot\mathbf{u}_{j}=\frac{\pi}{2}{\hbar\omega}\mathcal{J}_{ij}^{(S)}(\omega),$ (100)
@@ B.5 Mutual Poynting theorem
$\mathscr{P}_{ij}^{abs}=\frac{\omega}{2}\varepsilon_{0}\int_{V}d^{3}\mathbf{r}\,\text{Im}[\varepsilon_{\omega}(\mathbf{r})]\mathbf{E}_{i}(\mathbf{r})\cdot\mathbf{E}_{j}^{*}(\mathbf{r})\\ =-\oint_{\partial\Omega}d^{2}\mathbf{r}\,\mathbf{S}_{ij}\,\cdot\mathbf{m},$ (101)
@@ Appendix C Correlators
In this Section we compute the elements $C_{ij}^{(M)}(\tau)$ of the correlator matrix of the medium-assisted reservoir; a similar result holds for the elements of the correlator matrix of the scattering-assisted reservoir.
We consider the following scenario: (a) the initial quantum state of the electromagnetic environment is a product state, that is, $\hat{\rho}_{E}(0)=\hat{\rho}_{E}^{(M)}(0)\otimes\hat{\rho}_{E}^{(S)}(0)$, where $\hat{\rho}_{E}^{(M)}(0)$ and $\hat{\rho}_{E}^{(S)}(0)$ are the initial density operators of the medium-assisted reservoir and the scattering-assisted reservoir, respectively; (b) the two reservoirs are initially in thermal states at temperatures $T_{0}^{(M)}$ and $T_{0}^{(S)}$.
The correlation function $C^{(M)}_{ij}(\tau)$ is given by
$C^{(M)}_{ij}(\tau)=\langle\hat{F}^{(M)}_{i}(\tau)\hat{F}^{(M)}_{j}(0)\rangle$ (102)
where $\hat{{F}}^{(M)}_{i}(\tau)$ is the interaction operator of the medium-assisted reservoir in the interaction picture,
$\hat{{F}}^{(M)}_{i}(\tau)=\hat{U}^{\dagger}_{E}(\tau)\hat{F}^{(M)}_{i}\hat{U}_{E}(\tau),$ (103)
$\hat{U}_{E}(\tau)=\exp(-i\hat{H}_{E}^{\text{bright}}\tau/\hbar)$ is the free evolution operator of the bright bosonic modes of the electromagnetic environment, and $\langle\cdot\rangle=\text{Tr}_{E_{M}}[\cdot\,\rho_{E}^{(M)}(0)]$. Starting with 103, and using Eqs. 32, 46, and the commutation relations 30 we obtain
$\hat{F}_{i}^{(M)}(\tau)=\sum_{j=1}^{N}\int_{0}^{\infty}d\omega[G_{ij}^{(M)}(\omega)\hat{C}_{j}(\omega)e^{-i\omega\tau}+h.c.].$ (104)
Substituting 104 into 102, and using again 30 we find
$C_{ij}^{(M)}(\tau)=\int_{0}^{\infty}d\omega\sum_{n=1}^{N}\left[G^{(M)}_{in}(\omega)G^{(M)*}_{jn}(\omega)\langle\hat{C}_{n}(\omega)\hat{C}^{\dagger}_{n}(\omega)\rangle e^{-i\omega\tau}+G^{(M)*}_{in}(\omega)G^{(M)}_{jn}(\omega)\langle\hat{C}^{\dagger}_{n}(\omega)\hat{C}_{n}(\omega)\rangle e^{+i\omega\tau}\right],$ (105)
@@
$C_{ij}^{(M)}(\tau)=\\ \int_{0}^{\infty}\mathrm{d}\omega[(1+n_{\omega}^{(M)}){M}_{ij}(\omega)e^{-i\omega\tau}+n_{\omega}^{(M)}{M}_{ij}^{*}(\omega)e^{i\omega\tau}].$ (108)
@@ Appendix D Surface Integral Equation Method
Following Poggio–Miller–Chang–Harrington–Wu–Tsai (PMCHWT) [69, 70, 71, 72], the equivalent electric and magnetic surface currents $\mathbf{j}_{e}(\mathbf{r})$ and $\mathbf{j}_{m}(\mathbf{r})$, defined on $\partial V$, solve:
$\boldsymbol{\mathcal{Z}}\,\mathbf{j}=\mathbf{v},$ (109)
@@
$\mathbf{j}=\left[\mathbf{j}_{e},\mathbf{j}_{m}\right]^{T}$, $\mathbf{v}=\left[\mathbf{e}_{0},\mathbf{h}_{0}\right]^{T}$, $\mathbf{e}_{0}=\left.-\mathbf{n}\times\mathbf{n}\times{\bf{E}}_{inc}\right|_{\partial V}$, and $\mathbf{h}_{0}=\left.-\mathbf{n}\times\mathbf{n}\times{\bf{H}}_{inc}\right|_{\partial V}.$ The EFIE/MFIE boundary operators $\mathcal{T}_{\pm}$ and $\mathcal{K}_{\pm}$ act on a tangential test function $\mathbf{w}$ as:
$\boldsymbol{\mathcal{K}}_{\pm}$ $\left\{{\bf w}\right\}\left({\bf r}\right)=\mathbf{n}\times\mathbf{n}\times\int_{\partial V}{\bf w}\left({\bf r}^{\prime}\right)\times\nabla^{\prime}g^{\pm}\left({\bf r}-{\bf r}^{\prime}\right)dS^{\prime},$ (111a)
$\boldsymbol{\mathcal{T}}_{\pm}$ $\left\{{\bf w}\right\}\left({\bf r}\right)=-ik^{\pm}_{\omega}\mathbf{n}\times\mathbf{n}\times\int_{\partial V}g^{\pm}\left(\mathbf{r}-\mathbf{r}^{\prime}\right){\bf w}\left({\bf r}^{\prime}\right)dS^{\prime}$
$-\frac{1}{ik^{\pm}_{\omega}}\mathbf{n}\times\mathbf{n}\times\int_{\partial V}\nabla^{\prime}g^{\pm}\left(\mathbf{r}-\mathbf{r}^{\prime}\right)\nabla^{\prime}_{S}\cdot{\bf w}\left({\bf r}^{\prime}\right)dS^{\prime},$ (111b)